The Brenmar Sales Company had a gross profit margin (gross profits divided by sales) of 29 percent and sales of $8.1 million last year. 78 percent of the firm's sales are on credit, and the remainder are cash sales. Brenmar's current assets equal $1.8 million, its current liabilities equal $298,800 and it has $107,300 in cash plus marketable securties.

A. If Brenmar's accounts receivale equal $561,600, what is is average collection period?

B. If Brenmar reduces its average collection period to 15 days, what will be its new level of accounts receivable?

C. Brenmar's inventory turnover ratio is 8.9 times. What is the level of Brenmar's inventories?
Answer #1

a.Credit Sales =78%*8100000=6318000
Average Collection Period =365/(Credit Sales/Account Receivables) =365/(6318000/561600) =32.44

b. Average Collection Period=365*Account Receivables/Credit Sales
New Account Receivables =Average Collection Period*Credit Sales/365 =15*6318000/365=259643.84

c. Gross Profit =Sales*Gross Profit Margin =8,100,000*29% =2349000
COGS =Sales-Gross Profit =8100000-2349000 =5751000
Inventory =COGS/Inventory Turnover Ratio=5751000/8.9 =646179.78